OpenStudy (anonymous):

Is the answer to this C? Consider the function y=-2-3cos(x+π). What effect does the -2 have on the basic graph? A. Horizontal shift right 2 units B. Vertical stretch by factor 2 C. Vertical shift down 2 units D. Horizontal shift left 2 units

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Okay wait y = -3(cos(pi + x)) -2 y = 3cosx -2 Now when x = 0 y = 3-2 =1

OpenStudy (anonymous):

Good nothing bad ...C ...Yea C Sorry I was confirming my thoughts